OpenStudy (anonymous):

Solve using the substitution method: x+3y=10 xy=7

OpenStudy (anonymous):

HI JUANITA I DID NOT SEE UR REPLY.....FIND THE LCM OF 6Y^4 AND 18Y^5

hero (hero):

x + 3y = 10 x = 10 - 3y x = 7/y x = x 10 - 3y = 7/y y(10-3y) = 7y 10y - 3y^2 = 7y 3y = 3y^2 3y^2 - 3y = 0 3y(y - 1) = 0 y = 0, 1

OpenStudy (anonymous):

I typed in that answer but it was incorrect. I need two points that will equal both equations.

hero (hero):

Notice, I only solved for y, not x

hero (hero):

Clearly that is only values for y

hero (hero):

you mean you need one point with two coordinates (x,y)

hero (hero):

By the way the point is (1,7)

OpenStudy (anonymous):

So would I plug y into the equation to get x?

hero (hero):

Yes, correct. We had two values for y we could choose from. Either y = 0 or y = 1 would work. In this case, y = 0 is not a solution because xy = 7 => x*0 = 7 => 0 = 7 is not true

hero (hero):

So only y = 1 would work

OpenStudy (anonymous):

Ahhh okay! I understand now. Thank you!

hero (hero):

I think Luis Rivera posted another solution but I'm convinced that his is not correct.

hero (hero):

He has included another point

hero (hero):

x + 3y = 10 x = 10 - 3y x = 7/y x = x 10 - 3y = 7/y y(10-3y) = 7 10y - 3y^2 = 7 3y^2 - 10y + 7 = 0 3y^2-3y-7y+7 = 0 3y(y-1)-7(y-1)=0 (y-1)(3y-7)=0 y = 1 y = 3/7 x = 7 x = 3